# This manuscript is a preprint and has not been peer reviewed | Version: July 2025

### v1

@@ 3.5 Efficiency Analysis
In addition to accuracy, The computational efficiency of the replay based methods was measured using the total amount of replay batches performed during training. Additionally, For TFC-SR, the total number of Memory checks performed was also measured. The results for the primary experiments (buffer_capacity = 1000) and the “spaced replay” variant of TFC-SR are summarised in Table 3.
For the main experiments, both TFC-SR and Standard ER performed an identical number of replay batches (15800) with the only difference being the number of memory checks performed by TFC-SR (50). This indicates that the performance advantage of TFC-SR (13.17% vs 7.40%) stems from the stabilizing effect of its active recall probes, rather than the replay volume.
The trade-off between replay volume and accuracy was further explored in Appendix C, where a variant of TFC-SR with spaced replay achieved competitive performance of 10.18% - surpassing Standard ER - with over 77% fewer replay batches. This demonstrates the potential for significant computational savings when an adaptive schedule is allowed to reduce replay frequency.
Table 3: Efficiency comparison of replay based methods on Split CIFAR-100 benchmark. All methods use a buffer capacity of 1000.
@@ 4 Discussion
As detailed in Section 3, TFC-SR was compared against other prominent Continual Learning methods on two benchmarks - Split MNIST and Split CIFAR-100. On both benchmarks, it demonstrated strong performance and even significantly outperformed all other methods on the difficult Split CIFAR-100 benchmark. This demonstrated that TFC-SR is a highly effective strategy for continual learning. The results from the ablation studies further revealed a key trade-off, showing that TFC-SR’s advantage is more pronounced in memory-constrained settings.
@@
The performance gain of TFC-SR was achieved despite both methods performing an identical number of replay batches and the only difference being the presence of the active recall mechanism in TFC-SR. This suggests that the benefit stems not from the volume of replay, but from the act of memory check itself. We frame this mechanism as “Active Recall Probe”. We posit two primary reasons for the stabilizing effect. First, the repeated forward passes on past experiences force the network’s Batch Normalization layers to maintain their running statistics for older distributions, directly countering a known source of representational drift. Second, the periodic activation of neural pathways associated with past memories may act as a form of implicit regularization, strengthening those representations against interference from new learning. Additionally, This mechanism is incredibly efficient requiring minimal computational resources. As mentioned in section 3.5, It only required 50 probes for the entire Split CIFAR-100 benchmark.
@@
TFC-SR contributes a novel mechanism to the family of replay-based continual learning methods. Unlike premier regularization methods such as EWC [6] and SI [7], which indirectly protect old knowledge by penalizing weight changes, TFC-SR directly reinforces past representations through its Active Recall Probe. Our approach also differs from other advanced replay methods. While excellent methods like iCaRL [12] focus on intelligent exemplar management for replay buffer and interesting generative methods like those in Van De Ven et al., 2020 [13] and Tadros et al., 2022 [14] focus on creating replay data without storage, TFC-SR introduces a new dimension: the process of active, periodic memory retrieval as a learning mechanism in its own right.
@@
However, we acknowledge that this study is not without its own limitations. First, our experiments were conducted exclusively on image classification benchmarks; the efficacy of TFC-SR on other data modalities such as Natural Language Processing or Reinforcement Learning tasks remain to be explored. Second, our replay buffer used a standard reservoir sampling strategy. Incorporating more advanced buffer management techniques, such as iCaRL’s herding algorithm, Iterative Projection and Matching (IPM) [15] or gradient-based sample selection [16] could provide further improvements. Lastly, We used a naive notion of “mastery” for the adaptive scheduler. It was based on a simple performance threshold. A more nuanced metric could unlock more complex and efficient scheduling dynamics.
@@
The findings from this paper open several exciting avenues for future research. Our preliminary results with a ’Spaced Replay’ variant (Appendix C) show a promising trade-off between accuracy and computational cost, warranting a deeper investigation into sparse, episodic consolidation. Furthermore, the viability of an alternative ’Mastery-Gated Progression’ paradigm (Appendix D) suggests a rich area for future research into self-directed and adaptive learning schedules. Ultimately, this work strongly supports the hypothesis that building computational analogues of human cognitive processes, like active recall, is a fruitful direction for developing more robust, efficient, and capable artificial intelligence systems.
@@ Acknowledgments
I would like to offer my genuine gratitude to all the people who have provided their direct or indirect support throughout the completion of this work. I also acknowledge the significant contribution of AI assistance in improving readability and grammatical structure of this paper as well as in saving significant setup time during experimentation. I also would like to extend my thanks to my family and friends for their unwavering support and encouragement.
@@ Appendix A TFC-SR’s Sensitivity to Mastery Threshold
To find the optimal mastery threshold for TFC-SR, we tested its performance across a range of values on the CIFAR-100 benchmark: (10.0, 20.0, 30.0, 50.0, 70.0, 90.0, 99.0). We found that performance was robust across all values. The results are outlined in Table A.1 and illustrated in Figure A.1.
For thresholds between 10% and 90%, the final accuracy was nearly identical (13.17%), indicating that the model’s performance on the replay buffer consistently exceeded the targets. This caused the adaptive scheduler to adopt its most efficient mode, performing the minimum number of memory checks (50-63).
The 99% threshold served as a crucial stress test, creating a condition where the memory check would only succeed if the model’s retention on past knowledge was nearly perfect. The scheduler performed 103 memory checks in this configuration, significantly more than in the lower-threshold runs. This confirms that the check was frequently failing, forcing more intensive probing as designed.
However, the fact that only 103 checks were performed, which is close to only half of the maximum possible, proves that the model did succeed in meeting the 99% threshold on multiple occasions. This indicates moments of powerful, near-perfect memory consolidation even while learning new tasks.
Most notably, the final accuracy of this 99% threshold run was 11.58%. This result significantly outperforms the Standard ER baseline (7.40%), demonstrating the powerful benefit of Active Recall Probe mechanism itself. Additionally, this performance is only slightly lower than the performance of the best-performing threshold values (13.17%), showing that the algorithm is robust and does not catastrophically fail even when subjected to a highly demanding and volatile replay schedule.
Table A.1: TFC-SR’s performance across multiple mastery thresholds on CIFAR-100 benchmark
@@
Figure A.1: Learning curves of TFC-SR with varying mastery thresholds on Split CIFAR-100. For comparison, the performance of Standard ER baseline (buffer capacity = 1000) is included. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis)
@@ Appendix B Full Learning Curves for Buffer Capacity Analysis
We evaluated both TFC-SR and Standard ER across four different buffer capacities: 100, 500, 1000, and 2000. The full, task-by-task learning trajectory for each TFC-SR run is shown in Figure B.1. The corresponding trajectories for Standard ER baseline are shown in Figure B.2. The final accuracy values from these curves were used to generate the trend plot in Figure 3 and the summary in Table 2 of the main paper.
Figure B.1: Learning curves of TFC-SR with varying buffer capacities on Split CIFAR-100. Each line represents a different buffer capacity (100, 500, 1000, 2000). The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis)
Figure B.2: Learning curves of Standard ER with varying buffer capacities on Split CIFAR-100. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis)
@@ Appendix C Analysis of a Spaced Replay Variant (TFC-SR²)
To further investigate the trade-off between replay volume and final performance, we implemented and tested a “Spaced Replay” variant of TFC-SR, which we informally named TFC-SR2. In this setup, the continuous mixed batch replay is disabled; instead the replay is performed only during the specific epochs that are triggered by the adaptive scheduling mechanism. On all other epochs, the model trains exclusively on data from the new task. This experiment allows us to isolate the impact of the scheduled replay events themselves, under a drastically reduced computational budget. The results of this experiment are compared against our primary TFC-SR model and the Standard ER baseline in Table 3 of the main paper while the full learning curves are illustrated in Figure C.1.
The TFC-SR (Spaced Replay) variant achieved a final accuracy of 10.18%. This result is highly significant for two reasons. First, despite using over 77% fewer replay batches (3555 vs 15800), it substantially outperforms the brute-force Standard ER baseline (7.40%), demonstrating that even a small amount of intelligently scheduled replay is more effective than a high volume non-adaptive replay. Second, while its accuracy is lower than our main continuous replay variant (13.17%), this result confirms that a high volume of replay is still beneficial for achieving maximum performance on a challenging benchmark. This finding highlights a clear and practical trade-off between computational efficiency and performance.
Figure C.1: Learning curves comparing TFC-SR (Spaced Replay) against our primary TFC-SR model and the Standard ER baseline on Split CIFAR-100. All methods used a buffer size of 1000. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis).
@@ Appendix D Mastery-Gated Progression
In developing our approach, we explored a second, more literal interpretation of the “deliberate practice” analogy, which we named Mastery-Gated Progression (MGP). Unlike TFC-SR, which interleaves learning and replay, MGP separates them into two distinct phases. In this paradigm, the model trains on a new task until it achieves a performance greater than a predefined mastery threshold. Only after this criterion is met does the model “progress” to the next task. This creates an adaptive training schedule where the time spent on each task is an emergent property of the task’s difficulty.
The mastery criterion is two-fold: the model’s accuracy on the current task’s test set must exceed a new_task_mastery_thresh, and its performance on past tasks (measured via evaluate_replay_buffer) must remain above a retention_thresh. A max_epochs_per_task hyperparameter is used as a safeguard to prevent infinite training loops on exceptionally difficult tasks. The full logic is detailed in Algorithm D.1.
We tested MGP on the Split CIFAR-100 benchmark. Its performance is compared against the primary baselines in Figure D.1. The MGP model achieved a final accuracy of 11.63%, a result that is highly competitive with the Standard ER baseline (7.40%) and demonstrates the viability of this alternative learning paradigm.
The most insightful finding from this experiment was the variable number of epochs required to achieve mastery for each task: (47, 9,, 22, 50, 50, 9, 15, 18, 7, 19). This result makes the inherent difficulty of each incremental task explicit. For example, the model required 47 epochs to learn the first task from a random initialization, but only 9 for the second, suggesting a degree of positive knowledge transfer. This dynamic allocation of computational effort stands in contrast to fixed-epoch schedules and presents a rich area for future research into self-directed learning [17].
@@
Require: Replay Buffer b, task_id, Dataloader l
Require: Hyperparameters: max_epochs, mastery_threshold, retention_threshold
@@
// perform mixed batch training
@@
new_task_perf = evaluate(m. e.test_dataset)
retention_perf = evaluate_replay_buffer(m,, b)
if new_task_perf >= mastery_thresh and retention_perf >= retention_thresh then
@@
Figure D.1: Learning curves comparing MGP against TFC-SR and the Standard ER on Split CIFAR-100. All methods used a buffer size of 1000. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis).

### v2

@@ 3.5 Efficiency Analysis
In addition to accuracy, the computational efficiency of the replay-based methods on the Split CIFAR-100 benchmark was recorded using the total number of replay batches performed during training. For TFC-SR, the total number of memory checks performed was also measured. The results are summarized in Table 3 for the primary experiments (buffer_capacity = 1000) as well as for the “Spaced Replay” variant of TFC-SR.
In the main experiments, both TFC-SR and Standard ER performed an identical number of replay batches (15,800) with the only difference being the 50 memory checks performed by TFC-SR. This indicates that the performance advantage of TFC-SR (13.17% vs. 7.40%) stems from the stabilizing effect of its Active Recall Probe and adaptive scheduling, rather than the amount of replay volume alone.
The trade-off between replay volume and accuracy was further examined in Appendix C, where a variant of TFC-SR with Spaced Replay achieved competitive performance of 10.18%, surpassing Standard ER, while reducing the number of replay batches by over 77%. This highlights the potential for significant computational savings when an adaptive schedule is allowed to reduce the replay frequency.
Table 3: Efficiency comparison of replay-based methods on Split CIFAR-100 benchmark. All methods use a buffer capacity of 1000.
@@ 4 Discussion
As detailed in Section 3, TFC-SR was compared against other prominent continual learning methods on two benchmarks: Split MNIST and Split CIFAR-100. On both benchmarks, it demonstrated strong performance and significantly outperformed all other methods on the more challenging Split CIFAR-100 benchmark. These results indicate that TFC-SR is a highly effective strategy for continual learning. The results from the ablation studies further revealed a key trade-off, showing that TFC-SR’s advantage is more pronounced in memory-constrained settings.
@@
The performance gain of TFC-SR over Standard ER was achieved despite both methods performing an identical number of replay batches and the only difference being the presence of the active recall mechanism in TFC-SR. This suggests that the benefit stems not from the volume of replay, but from the act of memory checking itself. We refer to this mechanism as the “Active Recall Probe”. We posit two primary reasons for its stabilizing effect, whose contributions appear to depend on the model architecture. Our primary hypothesis is that repeated forward passes on past experiences compel the network’s Batch Normalization layers to maintain their running statistics for older distributions, directly countering a known source of representational drift. This is strongly supported by our results on the Split CIFAR-100 benchmark, where the ResNet-18 model (which contains Batch Normalization) showed a dramatic performance advantage for TFC-SR over Standard ER. Our secondary hypothesis is that periodic activation of neural pathways associated with past memories acts as a form of implicit regularization, strengthening those representations against interference. The results from the Split MNIST benchmark, which used a simple CNN without Batch Normalization layers, are consistent with this view. On this benchmark, TFC-SR’s advantage was minimal, suggesting that the implicit regularization effect alone provides a more subtle, though still present, benefit. Therefore, we believe that it is the synergy between this implicit regularization and the stabilization of Batch Normalization that explains TFC-SR’s significant advantage over the Standard Experience Replay baseline on the Split CIFAR-100 benchmark. Additionally, this mechanism is extremely efficient and requires minimal computational resources. As mentioned in section 3.5, it only required 50 probes for the entire Split CIFAR-100 benchmark.
@@
TFC-SR contributes a novel mechanism to the family of replay-based continual learning methods. Unlike premier regularization methods such as EWC (kirkpatrick_overcoming_2017, ) and SI (zenke_continual_2017, ), which indirectly protect old knowledge by penalizing weight changes, TFC-SR directly reinforces past representations through its Active Recall Probe. Our approach also differs from other advanced replay methods. While excellent methods like iCaRL (rebuffi_icarl_2017, ) focus on intelligent exemplar management strategies for replay buffer that prioritize informativeness and notable generative approaches like those in (van_de_ven_brain-inspired_2020, ) and (tadros_sleep-like_2022, ) focus on creating replay data without storage, TFC-SR introduces a new dimension: the process of active, periodic memory retrieval as a learning mechanism in its own right.
@@
Nonetheless, this study is not without its own limitations. First, our experiments were conducted exclusively on image classification benchmarks; the efficacy of TFC-SR on other data modalities such as natural language processing or reinforcement learning tasks remains to be explored. Second, our replay buffer used a standard reservoir sampling strategy. Incorporating more advanced buffer management techniques, such as iCaRL’s herding algorithm, Iterative Projection and Matching (IPM) (zaeemzadeh_iterative_2019, ) or gradient-based sample selection (aljundi_gradient_2019, ) could provide further improvements. Third, our adaptive scheduler relied on a naive notion of “mastery”. It was based on a simple performance threshold. A more nuanced metric could unlock more sophisticated and efficient scheduling dynamics. Lastly, this work was conducted independently without access to large-scale infrastructure. As a result, we focused on feasible settings with modest compute requirements. Future work could explore scaling TFC-SR to more diverse architectures and longer task sequences.
@@
The findings in this paper open several exciting avenues for future research. Our preliminary results with a "Spaced Replay" variant (Appendix C) show a promising trade-off between accuracy and computational cost, warranting a deeper investigation into sparse, episodic consolidation. Furthermore, our experiments with a performance-threshold-based progression baseline (Appendix D) — a simple, literal interpretation of deliberate practice — underscore the potential of adaptive task scheduling and self-directed learning strategies as fruitful directions for future work. Ultimately, our results support the hypothesis that building computational analogues of human cognitive processes, such as Active Recall, is a fruitful direction for the development of more robust, efficient, and capable artificial intelligence systems.
@@ Acknowledgments
We are genuinely grateful to all who have provided direct or indirect support throughout this work. We acknowledge the significant contribution of a Large Language Model (LLM) in the preparation of this manuscript. Specifically, the LLM was used to refine readability, correct grammatical structure, and polish writing. It also served as a tool for research ideation and brainstorming in the early stages of this project, and an LLM-based code assistant helped save significant time during the experimental phase. We thank our anonymous reviewers for their insightful feedback. We are also grateful to the inventor of tea, and to the various bugs in our code that were our companions for a significant portion of our journey—but eventually had to part ways. We extend our heartfelt thanks to our family and friends for their unwavering support and encouragement. Finally, we recognize that progress in continual learning is a collective effort, and we acknowledge the value of prior works not only as baselines, but as part of a guiding path that shaped our own contributions.
@@ Appendix A TFC-SR’s Sensitivity to Mastery Threshold
To find the optimal mastery threshold for TFC-SR, we tested its performance across a range of values on the Split CIFAR-100 benchmark: (10.0, 20.0, 30.0, 50.0, 70.0, 90.0, 99.0). We found that the performance was robust across all values. The results are summarized in Table A.1 and Figure A.1.
For thresholds between 10% and 90%, the final accuracy was nearly identical (13.17%), indicating that the model’s performance on the replay buffer consistently exceeded the targets. This led the adaptive scheduler to adopt its most efficient mode by performing the minimum number of memory checks (50–63). The 70% threshold run yielded a slightly lower accuracy of 12.45% with 51 memory checks, likely due to stochastic variation from random sampling during training.
The 99% threshold served as a crucial stress test, creating a condition where the memory check would only succeed if the model’s retention of past knowledge was nearly perfect. The scheduler performed 103 memory checks in this configuration, significantly more than in the lower-threshold runs. This confirms that the check was frequently failing, thereby triggering the intended increase in probing.
However, the fact that only 103 checks were performed, roughly half of the maximum possible, proves that the model did succeed in meeting the 99% threshold on multiple occasions. This indicates instances of near-perfect memory consolidation even while learning new tasks.
In particular, the final accuracy of this 99% threshold run was 11.58%. This result significantly outperforms the Standard ER baseline (7.40%), highlighting the benefit of the Active Recall Probe mechanism. Additionally, this performance is only slightly lower than that of the best-performing threshold values (13.17%), showing that the algorithm is robust and does not catastrophically fail even when subjected to a highly demanding and volatile replay schedule.
Table A.1: TFC-SR’s performance across multiple mastery thresholds on CIFAR-100 benchmark.
@@
Figure A.1: Learning curves of TFC-SR with varying mastery thresholds on Split CIFAR-100. For comparison, the performance of Standard ER baseline (buffer capacity = 1000) is included. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learned (x-axis).
@@ Appendix B Full Learning Curves for Buffer Capacity Analysis
We evaluated both TFC-SR and the Standard ER baseline across four different buffer capacities: 100, 500, 1000, and 2000. The complete task-by-task learning trajectories for each TFC-SR run are shown in Figure B.1. The corresponding trajectories for the Standard ER baseline are shown in Figure B.2. The final accuracy values from these curves were used to generate the trend plot in Figure 3 and the summary in Table 2 of the main paper.
Figure B.1: Learning curves of TFC-SR with varying buffer capacities on Split CIFAR-100. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learned (x-axis).
Figure B.2: Learning curves of Standard ER with varying buffer capacities on Split CIFAR-100. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learned (x-axis).
@@ Appendix C Analysis of a Spaced Replay Variant (TFC-SR²)
To further investigate the trade-off between replay volume and final performance, we implemented and tested a “Spaced Replay” variant of TFC-SR, which we informally named TFC-SR2. In this setup, continuous mixed-batch replay is disabled; instead, the replay is performed only during the specific epochs that are triggered by the adaptive scheduling mechanism. In all other epochs, the model trains exclusively on data from the current task. This experiment allows us to isolate the impact of the scheduled replay events themselves under a drastically reduced computational budget. The results of this experiment are compared with our primary TFC-SR baseline and the Standard ER baseline in Table 3 of the main paper, while the full learning curves are illustrated in Figure C.1.
The TFC-SR (Spaced Replay) variant achieved a final accuracy of 10.18%. This result is significant for two reasons. First, despite using over 77% fewer replay batches (3,555 vs. 15,800), it substantially outperforms the brute-force Standard ER baseline (7.40%), demonstrating that even a small amount of intelligently scheduled replay is more effective than high-volume, non-adaptive replay. Second, while its accuracy is lower than our main continuous replay variant (13.17%), this result confirms that a higher replay volume remains beneficial for achieving maximum performance on a challenging benchmark. This finding highlights a clear and practical trade-off between computational efficiency and performance.
Figure C.1: Learning curves comparing TFC-SR (Spaced Replay) against our primary TFC-SR model and the Standard ER baseline on Split CIFAR-100. All methods used a buffer size of 1000. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learned (x-axis).
@@ Appendix D Mastery-Gated Progression
To explore a more literal interpretation of the deliberate practice analogy, we implemented a performance-gated, curriculum-style baseline, which we refer to as Mastery-Gated Progression (MGP) for convenience. While conceptually similar to existing curriculum and self-paced learning strategies, MGP is adapted to our continual learning setting. In contrast to TFC-SR’s interleaved training and replay, MGP separates them into distinct phases: the model trains on each task until it surpasses a predefined mastery threshold, at which point it progresses to the next task. This results in an adaptive schedule where training time on each task emerges from its inherent difficulty.
MGP uses a two-part mastery criterion: the model must achieve a test accuracy above a mastery_threshold on the current task, while maintaining performance above a retention_threshold on previously seen tasks. To prevent indefinite training on particularly difficult tasks, a max_epochs parameter is used as a safeguard. The full procedure is outlined in Algorithm D.1.
We tested MGP on the Split CIFAR-100 benchmark. Its performance is compared with the primary baselines in Figure D.1. The MGP model achieved a final average accuracy of 11.63%, outperforming the Standard ER baseline (7.40%). While not surpassing our main method (TFC-SR), this result suggests that performance-gated progression is a viable training strategy and merits further exploration.
The most insightful finding from this experiment was the variable number of epochs required to achieve mastery for each task: (47, 9, 22, 50, 50, 9, 15, 18, 7, 19). This result makes explicit the inherent difficulty of each incremental task. For example, the model required 47 epochs to learn the first task from a random initialization, but only 9 for the second, suggesting a degree of positive knowledge transfer. This dynamic allocation of training effort, driven by task difficulty and retention, contrasts with fixed-epoch schedules and highlights the potential for more adaptive, self-directed approaches [17] to continual learning.
@@
Require: Replay Buffer b, task_id
Require: Hyperparameters:
max_epochs, mastery_threshold, retention_threshold
@@
// perform mixed-batch training
@@
new_task_perf = evaluate(m, e.test_dataset)
retention_perf = evaluate_replay_buffer(m, b)
if new_task_perf >= mastery_threshold and
retention_perf >= retention_threshold then
@@
Figure D.1: Learning curves comparing MGP with TFC-SR and the Standard ER on Split CIFAR-100. All methods used a buffer size of 1000. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learned (x-axis).